How do I find the slope of f(0)=4, f(4)=0?
Answer by stanbon(75887)

How do I find the slope if f(0)=4, f(4)=0?
---
You have two points: (0,4) and (4,0)
-----
slope = (4-0)/(0-4) = -1
-----
(0,4) is the intercept
-----
So y = -x + 4
=================
Cheers,
Stan H.